OpenStudy (anonymous):

Last questions fo the school year :D What is the simplified form of the expression sqrt(16c^4) - sqrt(c^2) + 3sqrt(c^2) + sqrt(9c^2)

OpenStudy (anonymous):

\[\sqrt{16c^4} - \sqrt{c^2} + 3\sqrt{c^2} + \sqrt{9c^2}\]

OpenStudy (anonymous):

\[4c ^{2}+5c\]

OpenStudy (anonymous):

√16c^4 - √c² + 3√c² + √9c² = 4c² - c + 3c + 3c = 4c² +5c